The displacement of an object is given by x =sin2pi (t+0.01)m. Find amplitude of the oscillatory motion ,time period, maximum velocity , maximum acceleration and initial displacement

1
Expert's answer
2022-04-14T15:45:46-0400

"A=1~m,"

"T=\frac{2\pi}{\omega}=1~s,"

"v=\omega A=6.3~\frac m s,"

"a=\omega v=39.5~\frac m{s^2},"

"x_0=\sin 0.02\pi=0.06~m."